# Genesis of Axion Dark Matter via Black Hole and Cosmic String Dynamics

### v1

@@ 1 Introduction
Peccei-Quinn symmetry has been proposed in the Standard Model as a potential solution to the strong CP problem [1]. The pseudo-goldstone boson of this model, called the axion, [2] also serves as a dark matter candidate [3, 4, 5]. This $U_{PQ}(1)$ symmetry when spontaneously broken, leads to the formation of global strings called cosmic axion strings [6], [7]. Multiple studies have shown that global cosmic strings decay into radiation which can be massless goldstone bosons, and massive particles like axions [8], [9] and even gravitational waves (GW) [10, 11, 12].
Research on cosmic strings also involves their interaction with black holes, which lead to interesting effects. Strings when attached to a rotating black hole have been proposed to extract the black hole’s rotational energy [13], [14]. Further, string interactions with primordial black holes (PBHs) suggest the formation of a large black hole-cosmic string network [15]. Depending on their trajectories, strings near black holes may be chaotically captured or scattered [16], [17], while circular strings can exhibit vibrational and oscillatory behavior around a Schwarzschild black hole [18], [19]. Previous studies by the authors have also explored cosmic string–black hole systems in low-mass X-ray binaries, suggesting that the influence of strings on black hole spin and the system’s orbital period could produce observable signatures, providing potential evidence for the existence of cosmic strings [20, 21, 22]. Recent work has also examined the role of primordial black holes (PBHs) as a potential source of axionlike particles (ALPs) through Hawking radiation [23].

### v2

@@ 1 Introduction
Peccei-Quinn symmetry has been proposed in the Standard Model as a potential solution to the strong CP problem [1]. The pseudo-goldstone boson of this model, called the axion, [2] also serves as a dark matter candidate [3, 4, 5, 6, 7]. This $U_{PQ}(1)$ symmetry when spontaneously broken, leads to the formation of global strings called cosmic axion strings [8], [9]. Multiple studies have shown that global cosmic strings decay into radiation which can be massless goldstone bosons, and massive particles like axions [10], [11] and even gravitational waves (GW) [12, 13, 14].
Research on cosmic strings also involves their interaction with black holes, which lead to interesting effects. Strings when attached to a rotating black hole have been proposed to extract the black hole’s rotational energy [15], [16]. Further, string interactions with primordial black holes (PBHs) suggest the formation of a large black hole-cosmic string network [17]. Depending on their trajectories, strings near black holes may be chaotically captured or scattered [18], [19], while circular strings can exhibit vibrational and oscillatory behavior around a Schwarzschild black hole [20], [21]. Further, recent studies have derived the modified relativistic orbits for cosmic string - black hole systems [22]. Previous studies by the authors have also explored cosmic string–black hole systems in low-mass X-ray binaries, suggesting that the influence of strings on black hole spin and the system’s orbital period and accretion structure could produce observable signatures, providing potential evidence for the existence of cosmic strings [23, 24, 25]. Recent work has also examined the role of primordial black holes (PBHs) as a potential source of axionlike particles (ALPs) through Hawking radiation [26].